OpenStudy (anonymous):

The formula P=B/1+r is used to determine what amount of principal P should be invested for one year at simpleinterest rate r in order to have B dollars after a year. Solve the formula for r.

OpenStudy (anonymous):

If you mean P = B/(1+r), then r = P/B - 1.